OpenStudy (anonymous):

how can you determine whether the inverse of a function is itself a function? Is there examples of a function whose inverse is a function and a function whose inverse is not a function?

OpenStudy (hba):

f ( x ) = (x-2)/(x-1) if y = (x-2)/(x-1) yx-y= x - 2 yx-x= -2+y x(y-1)=y-2 x = (y-2)/(y-1) so g ( x ) the inverse function is also (x-2)/(x-1)

OpenStudy (anonymous):

@hba your a life saver ; thanks cutie

ganeshie8 (ganeshie8):